One student can mow a yard in 3 hours; another student, in 4 hours. About how long would it take if they work together? Please show me how you got the answer.
Answer by edjones(8007)

In work problems turn things upside down; add them; turn them right side up for the answer.
1st student: does 1/3 of the job in 1 hr
2nd student: does 1/4 fo the job in 1 hr
.
1/3+1/4=4/12+3/12=7/12 of the job gets done in 1 hr by both of them.
.
12/7 hrs it takes them to do it together or 1 hr 5/7 or about 1 hr 43 min
Ed